The school community service clud has 45 members. Of these 45 memders, 3\5 are boys. How many boys ar members of the Community ServiceClud?
Answer by texttutoring(324)

There are 45 total members, and 3/5 are boys. Just multiply 45 by 3/5:

45*3/5 = 135/5 = 27

27 of the members are boys.